Circular No. 6337

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 T. Dotani, Y. Ueda, M. Ishida, F. Nagase, and H. Inoue, Institute

of Space and Astronautical Science; and Y. Saitoh, University of Tokyo, communicate on behalf of the ASCA team: “We observed GRO J1744–28 with ASCA between Feb. 26.42 and 27.09 UT with a net exposure time of ∼ 20 000 s. We detected a bright bursting/pulsating source at a location consistent with the VLA position (IAUC 6323). The GIS intensity of the persistent component in the range 1–10 keV was ∼ 130 counts/s (after deadtime correction). The energy spectrum of the persistent emission is hard, is contaminated by a large soft x-ray absorption, and can be approximated by a power law with photon index α = 1.22 and absorbing column density N_{H} = 5.1 × 10^{22} cm^{−2}. In addition, a structure is seen that appears as a broad hump around 6 keV and an edge at 7 keV and that can be attributed to a highly absorbed component. Assuming a cosmic abundance of iron, one requires a column density of 2 × 10^{23} cm^{−2} to explain the 7keV edge structure. With the inclusion of this highly-absorbed component, the photon index of the power-law component becomes a little steeper, α ∼ 1.36 ± 0.02, with N_{H} = (5.3 ± 0.1) × 10^{22} cm^{−2}. The average flux of the persistent component was 1.9 × 10^{−8} erg s^{−1} cm^{−2} in the range 2–10 keV, which corresponds to ∼ 900 mCrab. The apparent barycentric pulse period was found to be 0.467033(5) s on Feb. 26.76 UT. The pulse profile is almost sinusoidal with a single peak, and the pulse fraction in the band 2–10 keV is ∼ 10 percent (peak-to-peak). After correction for the orbital doppler effect using the parameters on IAUC 6286, the intrinsic pulse period at this epoch is estimated to be 0.467039 s. A total of nine bursts was detected during ∼ 16 000 s of high-time-resolution data. The duration of bursts is ∼ 10 s, and their peak flux is an order of magnitude larger than the persistent flux. After the decay of the bursts, the persistent flux decreased by 30 percent and then returned to the preburst level within a few hundred seconds. The energy spectrum of bursts is similar to, but slightly harder than, the persistent emission. The burst spectrum can be well fitted by a power law with α = 1.01 ± 0.04 and N_{H} = (6.1 ± 0.2) × 10^{22} cm^{−2}.”

COMET C/1995 Y1 (HYAKUTAKE) Visual m_{1} estimates: Feb. 18.25 UT, 8.0 (A. Pereira, Cabo da Roca,

Portugal, 9×34 binoculars); 28.23, 8.1 (J. Shanklin, Minera, Wales, 14×100 binoculars); Mar. 2.77, 8.7 (A. Pearce, Lockington, Victoria, 20×80 binoc.).

1996 March 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green